# O1-Pruner: Length-Harmonizing Fine-Tuning for O1-Like Reasoning Pruning

### v1

@@ 5.2 Results
Across both models, Marco-o1-7B and QwQ-32B-Preview, O1-Pruner outperforms other methods in average length of generated solutions, with a noticeable improvement on accuracy. For instance, in the Marco-o1-7B experiments, O1-Pruner achieves an average accuracy of 76.8%, accompanied by a 40.5% reduction in solution length compared to the baseline. Similarly, for QwQ-32B-Preview, O1-Pruner yields an average accuracy of 89.3%, with a 34.7% reduction in solution length. These improvements highlight the robustness of O1-Pruner in enhancing computational efficiency without sacrificing accuracy.
@@
The Fast-Solving Prompt method, while achieving a moderate reduction in solution length, compromises accuracy in most cases. This trade-off is evident from its lower AES values compared to O1-Pruner, indicating that the reduction in reasoning length often comes at the cost of problem-solving performance. On the other hand, SFT provides a better balance than the Fast-Solving Prompt, but its improvements in reasoning length remain marginal, with limited gains in AES.
The DPO method achieves a reasonable balance between accuracy and length, but it falls short of the performance achieved by O1-Pruner. Besides, the average accuracy decreases notably on Marco-o1-7B.
5.3 Inference Time-Cost
In this subsection, we take the MATH test set as an example to explore the time overhead during the model inference phase. We utilize one A800 GPU and the VLLM (Kwon et al., 2023) library for inference, recording the average inference time. For the Marco-o1 model, we employ one A800 GPU, while for the QwQ-32B-Preview model, we use four A800 GPUs. As illustrated in Fig. 3, the inference time results reveal notable differences across methods and models:
For the Marco-o1-7B model, the Baseline approach demonstrates an inference time of approximately 2 minutes, while the Fast-Solving Prompt and SFT methods achieve slightly shorter times. Both the DPO and O1-Pruner methods exhibit significantly reduced inference times, with O1-Pruner achieving the shortest duration, slightly exceeding 1 minute.
@@ 5.3 Inference Time-Cost
For the larger model QwQ-32B-Preview, the overall inference time is considerably higher. The Baseline approach records the longest inference time, approaching 6 minutes, while the DPO and SFT methods achieve slightly shorter durations. Notably, the Fast-Solving Prompt reduces the inference time to around 5 minutes, likely due to the strong instruction-following capabilities of large models. Once again, O1-Pruner demonstrates the shortest duration, achieving an inference time of approximately 4 minutes.
These results demonstrate that O1-Pruner can effectively accelerate the inference of long-thought models and is applicable to both smaller and larger language models.
@@
In summary, O1-Pruner represents a significant advancement in optimizing long-thought reasoning for math problem-solving tasks, achieving the best balance between accuracy and efficiency while minimizing computational overhead.
6 Further Evaluation
@@
Table 3: Ablation experiments on $\lambda$. Overall, the model’s accuracy and solution length increase with the penalty coefficient $\lambda$. $\lambda$ = 2 achieves an optimal balance between accuracy and efficiency.
6.1 Ablations
In this subsection, we conduct an ablation study on the accuracy constraint coefficient $\lambda$. We select several different values of $\lambda$ ($\lambda=0,1,2,5$) and evaluate the model accordingly. For the sake of brevity, we only report the average metrics across three datasets. It can be observed that, overall, the model’s accuracy increases as the penalty coefficient lambda rises, while the required inference length also grows. In our experiments, for Marco-o1-7b, setting $\lambda=2$ achieves a favorable trade-off between accuracy and efficiency.
@@ 6.1 Ablations
6.2 Performance Under Dataset of Different Difficulty Levels
In this subsection, we investigate the performance and characteristics of O1-Pruner across datasets of varying difficulty levels. Due to limited computational resources, we exclusively selected Marco-o1 for experimentation. Utilizing the data constructed from the MATH dataset as mentioned in prior experiments (comprising 5k problems * 16 solutions), we partition the dataset into three subsets of differing difficulty based on model’s average accuracy on them.
In Fig 4, We observe that models trained on more challenging datasets tend to generate longer solutions, as these datasets typically contain problems requiring more complex solutions. At the same time, by learning the correct solutions of harder problems, the models improve their problem-solving capabilities and ultimately achieve higher accuracy. In contrast, for the least challenging datasets, although the generated solution lengths are reduced, there is no improvement in accuracy. These experimental results suggest that while our approach demonstrates significant effectiveness in optimizing long-thought reasoning, it remains highly influenced by the nature of the training data.

### v2

@@ 5.2 Experimental Results
Table 2 demonstrates the performance of various methods across different evaluation metrics. The proposed O1-Pruner consistently achieves superior performance in balancing reasoning accuracy and efficiency compared to baseline and competing methods. Notably, it exhibits the best trade-off between accuracy and reasoning length across all datasets, as further supported by its significantly higher Accuracy-Efficiency Score (AES) values. Across both models, Marco-o1-7B and QwQ-32B-Preview, O1-Pruner outperforms other methods in average length of generated solutions, with a noticeable improvement on accuracy. For instance, in the Marco-o1-7B experiments, O1-Pruner achieves an average accuracy of 76.8%, accompanied by a 40.5% reduction in solution length compared to the baseline. Similarly, for QwQ-32B-Preview, O1-Pruner yields an average accuracy of 89.3%, with a 34.7% reduction in solution length. These improvements highlight the robustness of O1-Pruner in enhancing computational efficiency without sacrificing accuracy.
@@
The Fast-Solving Prompt method, while achieving a moderate reduction in solution length, compromises accuracy in most cases. This trade-off is evident from its lower AES values compared to O1-Pruner, indicating that the reduction in reasoning length often comes at the cost of problem-solving performance. On the other hand, SFT provides a better balance than the Fast-Solving Prompt, but its improvements in reasoning length remain marginal, with limited gains in AES. The DPO method achieves a reasonable balance between accuracy and length, but it falls short of the performance achieved by O1-Pruner. Besides, the average accuracy decreases notably on Marco-o1-7B.
5.3 Inference Time-Cost Analysis
@@
In this subsection, we take the MATH test set as an example to explore the time overhead during the model inference phase. We utilize one A800 GPU and the VLLM (Kwon et al., 2023) library for inference, recording the average inference time. For the Marco-o1 model, we employ one A800 GPU, while for the QwQ-32B-Preview model, we use four A800 GPUs. As illustrated in Figure 3, the inference time results reveal notable differences across methods and models: For the Marco-o1-7B model, the baseline approach demonstrates an inference time of approximately 2 minutes, while the Fast-Solving Prompt and SFT methods achieve slightly shorter times. Both the DPO and O1-Pruner methods exhibit significantly reduced inference times, with O1-Pruner achieving the shortest duration, slightly exceeding 1 minute. For the larger model QwQ-32B-Preview, the overall inference time is considerably higher. The Baseline approach records the longest inference time, approaching 6 minutes, while the DPO and SFT methods achieve slightly shorter durations. Notably, the Fast-Solving Prompt reduces the inference time to around 5 minutes, likely due to the strong instruction-following capabilities of large models. Once again, O1-Pruner demonstrates the shortest duration, achieving an inference time of approximately 4 minutes.
@@
In summary, O1-Pruner represents a significant advancement in optimizing long-thought reasoning for math problem-solving tasks for both smaller and larger language models, achieving the best balance between accuracy and efficiency while minimizing computational overhead.
Table 3: Ablation experiments on $\lambda$. Overall, the model’s accuracy and solution length increase with the penalty coefficient $\lambda$. A larger $\lambda$ implies that the model places greater emphasis on variations in accuracy, thereby partially weakening the optimization for sequence length. $\lambda$ = 2 achieves an optimal balance between accuracy and efficiency.
@@
5.4 Ablation Study
Ablation on Hyper-parameter Sensitivity. In this part, we evaluate the hyperparameter sensitively of constraint coefficient $\lambda$. We select several different values of $\lambda$ ($\lambda=0,1,2,5$) and evaluate the model accordingly. For the sake of brevity, we only report the average metrics across three datasets. It can be observed that, overall, the model’s accuracy increases as the penalty coefficient lambda rises, while the required inference length also grows. In our experiments, for Marco-o1-7b, setting $\lambda=2$ achieves a favorable trade-off between accuracy and efficiency.
@@ 5.4 Ablation Study
Ablation on Difficulty Levels. We investigate the performance and characteristics of O1-Pruner across datasets of varying difficulty levels. Due to limited computational resources, we exclusively selected Marco-o1 for experimentation. Utilizing the data constructed from the MATH dataset as mentioned in prior experiments (comprising 5k problems * 16 solutions), we partition the dataset into three subsets of differing difficulty based on the model’s average accuracy. In Figure 4, We observe that models trained on more challenging datasets tend to generate longer solutions, as these datasets typically contain problems requiring more complex solutions. At the same time, by learning the correct solutions of harder problems, the models improve their problem-solving capabilities and ultimately achieve higher accuracy. In contrast, for the least challenging datasets, although the generated solution lengths are reduced, there is no improvement in accuracy. These experimental results suggest that while our approach demonstrates significant effectiveness in optimizing long-thought reasoning, it remains highly influenced by the nature of the training data.